If the price of labor is $5 and the price of capital is $10, what is the marginal rate of technical substitution at the optimal input choice?

0.5

2

decreasing

increasing

none of the above
Answer #1

at optimal input choice , marginal rate of technical substitution = price of labor / price of capital

In microe conomic theory, the Marginal Rate of Technical Substitution —or Technical Rate of Substitution —is the amount by which the quantity of one input has to be reduced when one extra unit of another input is used, so that output remains constant.

MRTS = PL / Pk = 5 / 10 = 0.5